# Decision-aware training of spatiotemporal forecasting models to select a top K subset of sites for intervention

### v2

@@ Abstract
Optimal allocation of scarce resources is a common problem for decision makers faced with choosing a limited number of locations for intervention. Spatiotemporal prediction models could make such decisions data-driven. A recent performance metric called fraction of best possible reach (BPR) measures the impact of using a model’s recommended size K subset of sites compared to the best possible top-K in hindsight. We tackle two open problems related to BPR. First, we explore how to rank all sites numerically given a probabilistic model that predicts event counts jointly across sites. Ranking via the per-site mean is suboptimal for BPR. Instead, we offer a better ranking for BPR backed by decision theory. Second, we explore how to train a probabilistic model’s parameters to maximize BPR. Discrete selection of K sites implies all-zero parameter gradients which prevent standard gradient training. We overcome this barrier via advances in perturbed optimizers. We further suggest a training objective that combines likelihood with a decision-aware BPR constraint to deliver high-quality top-K rankings as well as good forecasts for all sites. We demonstrate our approach on two where-to-intervene applications: mitigating opioid-related fatal overdoses for public health and monitoring endangered wildlife.
Abstract
This supplement includes additional experimental results and information for understanding and reproducing experiments.
@@ 1 Introduction
Statistical machine learning methods for spatiotemporal forecasting can play a vital role in high-stakes applications from mitigating overdoses in public health (Marks et al., 2021b) to forest fire management (Cheng & Wang, 2008) to wildlife monitoring (Golden et al., 2022; Hefley et al., 2017). Across these domains, there is a pressing need for predictive models that can make accurate predictions of near-term future events at fine spatiotemporal resolutions. Such models can enable stakeholders to make data-driven decisions about how to allocate limited resources to maximize utility.
@@
In this work, we are interested in helping decision-makers select where to intervene. That is, given historical data for a fixed set of $S$ candidate spatial sites, recommend a specific subset of given size $K$ for some kind of action or intervention. We think of hyperparameter $K$ as setting the budget for interventions. In an ideal world, decision-makers could afford interventions in all $S$ sites. However, when resource constraints allow only $K$ sites to receive interventions, a decision selecting a specific $K$-of-$S$ subset is required. While such decisions may often be heuristic in practice, we hope to offer data-driven solutions.
@@
With this goal in mind, choosing a sensible performance metric is critical to assessing which models have real-world utility. Common metrics such as squared error or absolute error are not well matched to where-to-intervene decisions because they treat all sites equally. Recent work on overdose forecasting has suggested a metric termed the fraction of best possible reach, or BPR (Heuton et al., 2022, 2024). BPR measures a ratio of event counts, where the numerator sums over the model’s recommended $K$ sites while the denominator sums over the best possible $K$ selected in hindsight. This type of evaluation has been used in a preregistered trial (Marshall et al., 2022a) for assessing forecasts of opioid overdoses in Rhode Island, as well as a follow-up feasability study (Allen et al., 2023). BPR is applicable to many where-to-intervene problems beyond public health. footnotetext: Code: https://github.com/tufts-ml/decision-aware-topk/
@@
While some publications have reported BPR in evaluations, we suggest that a natural goal would be for this performance metric to inform two other key parts of data-driven decision-making: model-based ranking and model training. By ranking, we mean that given fixed model parameters, the knowledge of BPR as the metric of interest should impact the numerical score assigned to each site to determine the top $K$. By training, we mean how to update model parameters to achieve good BPR. This paper contributes new methods for solving both ranking and training problems when BPR is the preferred metric.
@@
Our work overcomes two key technical barriers. The first barrier is in ranking. Given a fixed probabilistic model, determining how to compute a per-site score that is sorted to find the top $K$ is not obvious. It may be tempting to use the model’s per-site mean, but decision theory suggests different loss functions usually have distinct corresponding best estimators  (Berger, 2013; Murphy, 2022). As a relatively new metric, the problem of how to assign a numerical ranking to sites for optimal BPR decision-making is currently open. We contribute a tractable ranking method that is provably best for a reasonable bound on BPR. We further show how the score function estimator (Kleijnen & Rubinstein, 1996; Mohamed et al., 2020) can be used to calculate gradients of this ranking with respect to parameters.
@@
The second barrier prevents training parameters to improve BPR. First-order gradient descent enables training of many prediction methods, yet gradients of BPR with respect to parameters are problematic. The selection of the recommended top $K$ sites is piece-wise constant function of model parameters. While small changes to parameters will induce small changes per-site scores, only changes large enough to move a site into or out of the top-K ranked sites will have an affect on BPR. Thus, gradients of BPR with respect to parameters will be zero almost everywhere, preventing gradient methods from ever moving beyond subpar initial parameters. To fix this, we leverage recent advances in perturbed optimizers  (Abernethy et al., 2016; Berthet et al., 2020) to yield effective and efficient gradient estimation for BPR. Our team explored this idea for optimizing BPR alone in earlier non-archival work (Heuton et al., 2023); this paper offers an expanded treatment with more accessible presentation. Yet optimizing BPR alone can lead to model predictions with far lower likelihood than conventional training, leading to concerns about overall model quality and generalization. To address this, we further pose a constrained optimization problem to maximize likelihood subject to a BPR quality constraint that delivers quality top-K recommendations and good forecasts for all sites.
@@
We ultimately contribute methods for how-to-rank and how-to-train when making where-to-intervene decisions. Using these tools, a variety of models can be directly optimized to make effective top $K$ site recommendations. We demonstrate these contributions first on synthetic data, where we reveal how off-the-shelf methods without our innovations can be suboptimal for decision-making. We further offer evaluations of two real applications: mitigating opioid-related fatal overdoses and monitoring endangered birds.
@@
Throughout this paper, a recurring takeaway message is that conventional training methods based on maximizing likelihood can yield suboptimal decision-making for BPR, especially when forecasting models are misspecified. In this vein, our work shares the same broad goals as direct loss minimization (Wei et al., 2021) and loss-calibrated methods (Lacoste–Julien et al., 2011). We are inspired by the way these works bring utilities (losses) and Bayesian decision theory to improve training of probabilistic methods. Others have extended loss-calibrated methods to neural nets (Cobb et al., 2018) and to continuous (not discrete) actions (Kuśmierczyk et al., 2019). Yet a tractable and scalable recipe specifically for BPR-like objectives that prioritize $K$-of-$S$ decision-making is not an immediate next step from this past work. We hope our contributions spark interest in where-to-intervene problems in the methodological community and also lead to effective deployments of data-driven top-K decision-making in public health and beyond.
@@ 2 Technical Background and Problem Setup
Notation. Mathematically, some of our notation follows Sander et al. (2023). Function TopKMask takes as input a vector $\bm{r}$ of length $S$ and an integer $K$. It produces a binary vector of length $S$ with exactly $K$ entries equal to 1. Each 1 entry corresponds to a value in the top $K$ largest entries of the input vector. The $s$-th entry of the output is
[…]
@@
Problem definition. We wish to probabilistically model events that occur across $S$ distinct spatial sites over time. At each site, indexed by $s$, we can observe a non-negative scalar count or value $y_{s}\geq 0$. We assume that larger $y_{s}$ corresponds to greater value in intervention at site $s$. In our public health applications, $y_{s}$ represents counts of fatal overdoses. In our later wildlife monitoring case study, $y_{s}$ counts how often a rare animal appears. At each time $t$, we stack all observations into a vector $\bm{y}_{t}\in\mathbb{R}^{S}$.
Denote our joint model for this vector as $p_{\bm{\phi}}(\bm{y}_{t})$, where model parameter vector $\bm{\phi}$ defines the density over r.v. $\bm{y}_{t}$. We assume that explicitly evaluating this pdf and sampling values of $\bm{y}_{t}$ are both feasible. Given these assumptions, our framework is quite flexible: the vector $\bm{\phi}$ could represent the weights of a neural network or the coefficients of a logistic regression or a Bayesian hierarchical model.
@@
Definition of BPR. At current time $t$, we evaluate a model $\phi$’s ability to select a top-$K$ subset of sites for intervention before time $t+1$. Let $\mathcal{R}$ be the model’s recommended subset of $K$ sites among all $S$ sites. For the rest of this section, let $\bm{y}\in\mathbb{R}^{S}$ be the vector of observations at the target time $t+1$ (we skip time subscripts on $\bm{y}$ to keep notation simple). The vector $\bm{y}$ is not available when the decision of $\mathcal{R}$ is made. Following Heuton et al. (2022), we define BPR as:
@@
Suppose we have a loss function $L(\bm{r},\bm{y})$ (not necessarily related to BPR) that produces a scalar value indicating the overall quality of taking an “action” $\bm{r}$ and then realizing outcome $\bm{y}$. Lower values of $L$ indicate better decisions. A natural framework for making decisions about actions (Murphy, 2022) is to minimize the expected loss:
@@
As a simplistic example, if the loss is defined as the sum of squared errors, $\mathcal{L}(\bm{r},\bm{y})=\sum_{s=1}^{S}(r_{s}-y_{s})^{2}$, the optimal action is provably the per-site mean: $r_{s}=\mathbb{E}_{p_{\bm{\phi}}}[y_{s}]$. Similarly, a loss that sums up absolute errors has its optimal action the per-site median (Schwertman et al., 1990; Balkus, 2024). We tackle defining an optimal ranking for BPR.
@@
This way of writing the loss with dot products of top-K binary vectors is equivalent to $-\text{BPR}(\textsc{TopKIDs}(\bm{r},K),\bm{y})$.
@@ 3 Methods for Ranking
Many solutions exist: any vector $\bm{r}^{*}$ that satisfies $\textsc{TopKMask}(\bm{r}^{*},K)=\textsc{TopKMask}(\mathbb{E}_{p_{\phi}}[\bm{y}],K)$ can be an argmin. One valid solution here is the per-site mean $\bm{\bar{r}}(\phi)=\mathbb{E}_{p_{\phi}}[\bm{y}]$. However, this solution is only optimal for a potentially quite loose bound on BPR.
Tighter bound suggests the ratio estimator for ranking. Instead of bounding with constant $U$, we can upper bound the denominator in Eq. (4) by summing over all $S$ terms instead of the top $K$: $\sum_{s\in\textsc{TopKIDs}(\bm{y})}y_{s}\leq\sum_{s=1}^{S}y_{s}=\bm{1}\cdot\bm{y}$. This bound is tighter when the sum of all entries is less than $K\cdot U$, which is typically true of sparse $\bm{y}$ vectors in our applications. Our how-to-rank problem then becomes
@@ 4 Methods for Training
We now consider various ways to train the parameters of our probabilistic model on a training set that covers $T$ distinct time periods indexed by $t$. At each time, we have a data vector $\bm{y}_{t}\in\mathbb{R}_{\geq 0}^{S}$. Our model family in general factorizes $p(\bm{y}_{1:T})=\prod_{t=1}^{T}p_{\phi}(\bm{y}_{t}|\bm{y}_{1:t-1})$. To ease notation throughout, we omit conditioning on past history or other exogeneous features. So $p_{\phi}(\bm{y}_{t})$ below should be read as equal to $p_{\phi}(\bm{y}_{t}|\bm{y}_{1:t-1})$.
@@ 4.1 Maximum likelihood (ML) estimation
If the model is well-specified and training set size $T$ is large enough, this is a reliable strategy to estimate $\phi$. We could then use the ranking methods from Sec. 3 for where-to-intervene decisions. However, popular wisdom reminds us that “all models are wrong” in some way for real-world data. As we will show in later experiments, fitting a mispecified model via ML estimation can produce $\phi$ that deliver suboptimal BPR, even using the optimal ranking for that $\phi$.
@@ 4.2 Direct loss minimization for BPR
Inspired by the broad goal of direct loss minimization (Wei et al., 2021), another approach would be to find parameters that minimize our BPR-specific decision making loss $L^{\text{BPR}}$. In this strategy, we seek $\phi$ values that minimize
@@
To train with modern gradient methods, we’d need to compute the gradient $\nabla_{\phi}\mathcal{J}=\sum_{t}\nabla_{\phi}\bm{r}_{t}\cdot\nabla_{r_{t}}L_{t}$. However, technical difficulties arise with each term in this chain rule expansion. Below, we propose practical estimators for each term that overcome these barriers.
Gradient $\nabla_{\phi}\bm{r}_{t}$. The difficulty here is differentiating through an expectation, especially when $\bm{y}$ is a discrete random variable (integer counts in our overdose applications). We use the score function trick (Mohamed et al., 2020), popularized by Ranganath et al. (2014) yet dating back decades (Kleijnen & Rubinstein, 1996), sometimes also called REINFORCE (Williams, 1992). We can draw $M$ samples $\bm{y}^{(m)}_{t}\sim p_{\phi}$, then compute
@@
A downside of this estimator is high variance. We mitigate this with large $M$ values, though future work may use control variates (Ranganath et al., 2014; Mohamed et al., 2020) or try complex estimators for gradients of discrete expectations (Maddison et al., 2017; Dimitriev & Zhou, 2021).
Gradient $\nabla_{r_{t}}L_{t}$. For losses $L$ defined in terms of TopKMask binary vectors, like BPR, it is difficult to compute useful gradients because this loss is flat almost everywhere with respect to the input rankings $\bm{r}$. To overcome this barrier, we leverage recent advances in perturbed optimization (Berthet et al., 2020), also referred to as stochastic smoothing (Abernethy et al., 2016). A recent computer vision method (Cordonnier et al., 2021) shows how these ideas enable selecting a top-K set of patches from a high-resolution image for downstream prediction. We adapt this top-K approach to spatiotemporal forecasting for intervention.
Concretely, Cordonnier et al. (2021) obtain tractable $J$-sample Monte Carlo estimates of both the top-K indicator vector $\bm{b}=\textsc{TopKMask}(\bm{r},K)$ and the Jacobian $\nabla_{\bm{r}}\bm{b}$ needed for backpropagation. First, we draw $J$ independent samples of a standard Gaussian noise vector of size $S$: $\bm{z}_{j}\sim\mathcal{N}(0,I_{S})$. Then, we compute
@@
The Jacobian $\nabla_{\bm{r}}\bm{\hat{b}}$ is an $S\times S$ matrix, where entry $j,k$ gives the scalar derivative $\frac{\partial\bm{b}_{j}}{\partial r_{k}}$. The conceptual justification for the Jacobian estimator comes from Abernethy et al. (2016) and Berthet et al. (2020). Noise level $\sigma>0$ is a hyperparameter setting the strength of stochastic smoothing; it must be carefully selected in practice.
@@
When models are misspecified, directly estimating $\phi$ to minimize $\mathcal{J}^{\text{BPR}}$ should yield better BPR than some $\phi^{\prime}$ fit via conventional loss $\mathcal{J}^{\text{NLL}}$. However, probabilistic forecasts of near-future outcomes $\bm{y}_{t+1}$ produced by BPR-trained $\phi$ have questionable utility. Nothing in the $\mathcal{J}^{\text{BPR}}$ objective makes $p_{\phi}$ accurate even for $\bm{y}_{1:T}$; only relative ranking matters.

### v3

# Decision-aware Training of Spatiotemporal Forecasting Models to Select a Top-K Subset of Sites for Intervention

@@ Abstract
Optimal allocation of scarce resources is a common problem for decision makers faced with choosing a limited number of locations for intervention. Spatiotemporal prediction models could make such decisions data-driven. A recent performance metric called fraction of best possible reach (BPR) measures the impact of using a model’s recommended size K subset of sites compared to the best possible top-K in hindsight. We tackle two open problems related to BPR. First, we explore how to rank all sites numerically given a probabilistic model that predicts event counts jointly across sites. Ranking via the per-site mean is suboptimal for BPR. Instead, we offer a better ranking for BPR backed by decision theory. Second, we explore how to train a probabilistic model’s parameters to maximize BPR. Discrete selection of K sites implies all-zero parameter gradients which prevent standard gradient training. We overcome this barrier via advances in perturbed optimizers. We further suggest a training objective that combines likelihood with a BPR constraint to deliver high-quality top-K rankings as well as good forecasts for all sites. We demonstrate our approach on two where-to-intervene applications: mitigating opioid-related fatal overdoses for public health and monitoring endangered wildlife.
@@ 1 Introduction
Statistical machine learning methods for spatiotemporal forecasting can play a vital role in high-stakes applications from mitigating overdoses in public health (marksMethodologicalApproachesPrediction2021) to forest fire management (cheng2008ForestFire) to wildlife monitoring (golden2022SpatialPredictionWhooping; hefley2017DynamicSpatiotemporalWastingDisease). Across these domains, there is a pressing need for predictive models that can make accurate predictions of near-term future events at fine spatiotemporal resolutions. Such models can enable inform data-driven decisions about how to allocate limited resources to maximize utility.
@@
In this work, we seek to help decision-makers select where to intervene. Given historical data for a fixed set of $S$ candidate spatial sites, we develop models that can recommend a specific subset of given size $K$ for some action or intervention. We think of hyperparameter $K$ as setting the budget for interventions. In an ideal world, decision-makers could afford interventions in all $S$ sites. However, when resource constraints allow only $K$ sites to receive interventions, a decision selecting a specific $K$-of-$S$ subset is required. While such decisions may often be heuristic in current practice, we hope to offer data-driven solutions.
@@
With this goal in mind, choosing a sensible performance metric is critical to assessing which models have real-world utility. Common metrics such as squared error or absolute error are not well matched to where-to-intervene decisions because they treat all sites equally. Recent work on overdose forecasting has suggested a metric termed the fraction of best possible reach, or BPR (heutonPredictingSpatiotemporalCounts2022; heutonSpatiotemporalForecastingOpioidrelated2024). BPR measures a ratio of event counts. The numerator sums over the model’s recommended $K$ sites, while the denominator sums over the best possible $K$ selected in hindsight. This type of evaluation has been used in a preregistered trial (marshallPreventingOverdoseUsing2022) for assessing forecasts of opioid overdoses in Rhode Island, as well as a follow-up feasability study (allenTranslatingPredictiveAnalytics2023). BPR is applicable to many where-to-intervene problems beyond public health. footnotetext: Code: https://github.com/tufts-ml/decision-aware-topk
@@
While some publications have reported BPR in evaluations, we suggest that a natural goal would be for this performance metric to inform two other key parts of data-driven decision-making: model-based ranking and model training. By ranking, we mean that given fixed model parameters, the knowledge of BPR as the metric of interest should impact the numerical score assigned to each site to determine the top $K$. By training, we mean how to update model parameters to achieve high BPR. This paper contributes new methods for solving both ranking and training problems when BPR is the preferred metric.
@@
Our work overcomes several technical barriers. The first barrier is in ranking. Given a fixed probabilistic model, determining how to compute a per-site score, which will later be sorted to find the top $K$, is not obvious. It may be tempting to use the model’s per-site mean, but decision theory suggests not all loss functions recommend the per-site mean as the best estimator (berger2013StatisticalDecisionTheory; murphy2022Sec5_1DecisionTheory). As a relatively new metric, the problem of how to assign a numerical ranking to sites for optimal BPR decision-making is currently open. We contribute a tractable ranking method that is provably best for a reasonable bound on BPR. We further show how the score function estimator (kleijnen1996OptimizationScoreFunction; mohamed2020MonteCarloGradient) can be used to calculate gradients of this ranking with respect to parameters.
@@
The second barrier prevents training parameters to improve BPR. First-order gradient descent is a common, effective algorithm we would like to use. However, gradients of BPR with respect to parameters are problematic. While small changes to parameters induce some changes to per-site scores, only changes large enough to move a site into or out of the top-K ranked sites will adjust BPR. Thus, gradients of BPR with respect to parameters will be zero almost everywhere, preventing gradient methods from ever moving beyond subpar initial parameters. To fix this, we leverage recent advances in perturbed optimizers  (abernethyPerturbationTechniquesOnline2016; berthetLearningDifferentiablePerturbed2020) to yield effective and efficient gradient estimation for BPR. Our team explored this idea for optimizing BPR alone in earlier non-archival work (heutonlearning_dlm); this paper offers an expanded treatment with more accessible presentation, while addressing two more barriers.
The final barrier is designing a training objective to achieve applied goals. We find that optimizing BPR alone can lead to predictions with far lower likelihood than conventional training. This raises concerns about overall model quality and generalization. To address this, we pose a constrained optimization problem to maximize likelihood subject to a BPR quality constraint. This combined objective delivers quality top-K recommendations and good forecasts for all sites. Our objective is reminiscent of past additive combinations of a regression loss and decision loss (kao_directed_2009). Unlike that work, ours pursues non-convex losses and directly enforces decision quality via constraints.
@@
We ultimately contribute methods for how-to-rank and how-to-train when making where-to-intervene decisions. Using these tools, a variety of models can be directly optimized to make effective top $K$ site recommendations. We demonstrate these contributions first on synthetic data, where we reveal how off-the-shelf methods without our innovations can be suboptimal for decision-making. We further evaluate against alternatives on two applications: mitigating opioid-related fatal overdoses and monitoring endangered birds. We hope our contributions spark interest in where-to-intervene problems in the methodological community and also lead to effective deployments of data-driven top-K decision-making in public health and beyond.
Related work. The application of machine learning to decision-making problems is widely studied in operations research literature (bertsimasPredictivePrescriptiveAnalytics2020a; sadana_survey_2025), including the problem of how to train a model for downstream decision-making (mandi_decision-focused_2024). We review several model training approaches later in Sec. 4.
Other researchers have used decision-aware objectives to solve limited resource allocation problems. chung_decision-aware_2022 study how to allocate essential medicines in Sierra Leone across hospital sites. gupta_decision-aware_2024 pursue a where-to-intervene task in urban planning, selecting where to build speed humps to reduce pedestrian injuries. Our work differs in its focus on the BPR metric, our hybrid decision-aware objective that preserves likelihood, and evaluations that forecast the future given the recent past.
@@
Throughout this paper, a recurring takeaway is that conventional training based on maximizing likelihood can yield suboptimal decision-making for BPR, especially when forecasting models are misspecified. In this vein, our work shares similar goals as direct loss minimization (weiDirectLoss2021) and loss-calibrated methods (lacoste-julienApproximateInferenceLosscalibrated2011). We are inspired by the way these works combine task-specific losses and decision theory to improve probabilistic models. Others have extended loss-calibration to neural nets (cobbLossCalibratedApproximateInference2018a) and to continuous actions (kusmierczykVariationalBayesianDecisionmaking2019). Yet a tractable and scalable recipe that prioritizes top-$K$ where-to-intervene decision-making is not an immediate next step from this work.
@@ 2 Technical Background and Problem Setup
Notation. Mathematically, some of our notation follows sander_fast_2023. Function TopKMask takes as input a vector $\bm{r}$ of length $S$ and an integer $K$. It produces a binary vector of length $S$ with exactly $K$ entries equal to 1. Each 1 entry corresponds to a value in the top $K$ largest entries of the input vector. The $s$-th entry of the output is
$\textsc{TopKMask}(\bm{r},K)_{s}$ $=\begin{cases}1&\text{if }\textsc{rank}(\bm{r})_{s}\leq K\\ 0&\text{otherwise}\end{cases},$ (1)
@@
Problem definition. We wish to probabilistically model events that occur across $S$ distinct spatial sites over time. At each site, indexed by $s$, we can observe a non-negative scalar count or value $y_{s}\geq 0$. We assume that larger $y_{s}$ corresponds to greater value in intervention at site $s$. In our public health applications, $y_{s}$ represents counts of fatal opioid-related overdoses. In our later wildlife monitoring case study, $y_{s}$ counts how often a rare animal appears. At each time $t$, we stack all observations into a vector $\bm{y}_{t}\in\mathbb{R}^{S}$. We assume that the true data-generating distribution for each vector $\bm{y}_{t}$ given past history does not change over time, including between the training and test periods.
We denote our joint model for this vector as $p_{\bm{\phi}}(\bm{y}_{t})$, where model parameter vector $\bm{\phi}$ defines the density over r.v. $\bm{y}_{t}$. We assume that explicitly evaluating this pdf and sampling values of $\bm{y}_{t}$ are both feasible. Given these assumptions, our framework is quite flexible: the vector $\bm{\phi}$ could represent the weights of a neural network or the coefficients of a logistic regression or a Bayesian hierarchical model.
Given a training set of $T$ times, our model family in general factorizes $p(\bm{y}_{1:T})=\prod_{t=1}^{T}p_{\phi}(\bm{y}_{t}|\bm{y}_{1:t-1})$. To ease notation throughout, we omit conditioning on past history or other exogenous features. So $p_{\phi}(\bm{y}_{t})$ below should be read as equal to $p_{\phi}(\bm{y}_{t}|\bm{y}_{1:t-1})$ for models with such dependencies.
@@
Definition of BPR. At current time $t$, we evaluate a model $\phi$’s ability to select a top-$K$ subset of sites for intervention before time $t+1$. Let $\mathcal{R}$ be the model’s recommended subset of $K$ sites among all $S$ sites. For the rest of this section, let $\bm{y}\in\mathbb{R}^{S}$ be the vector of observations at the target time $t+1$ (we skip time subscripts on $\bm{y}$ to keep notation simple). The vector $\bm{y}$ is not available when the decision of $\mathcal{R}$ is made. Following heutonPredictingSpatiotemporalCounts2022, we define BPR as:
@@
Suppose we have a loss function $L(\bm{r},\bm{y})$ (not necessarily related to BPR) that produces a scalar value indicating the overall quality of taking an “action” $\bm{r}$ and then realizing outcome $\bm{y}$. Lower values of $L$ indicate better decisions. A natural framework for making decisions about actions (murphy2022Sec5_1DecisionTheory) is to minimize the expected loss:
@@
As a simplistic example, if the loss is defined as the sum of squared errors, $\mathcal{L}(\bm{r},\bm{y})=\sum_{s=1}^{S}(r_{s}-y_{s})^{2}$, the optimal action is provably the per-site mean: $r_{s}=\mathbb{E}_{p_{\bm{\phi}}}[y_{s}]$. Similarly, for the loss that sums up absolute errors, the optimal action is the per-site median (schwertman1990MedianProof; proofMedianMinimizesMAE). We tackle defining an optimal ranking for BPR.
@@
This way of writing the loss with dot products of top-K binary vectors is equivalent to $\text{BPR}(\textsc{TopKIDs}(\bm{r},K),\bm{y})$.
Connection to 0-1 knapsack. Given a fixed $\bm{y}$ vector, the problem of selecting $K$ sites to minimize $L^{\text{BPR}}$ can reduce to the canonical 0-1 knapsack problem (dantzigDiscreteVariableExtremumProblems1957) where each site $s$ would have value $y_{s}$ and weight 1, and the budget constraint allows just $K$ of all $S$ sites. Our how-to-rank contribution solves a more general problem: how to set $\bm{r}$ when $\bm{y}$ is not given but must be forecasted by our model.
@@ 3 Methods for Ranking
Many solutions exist: any vector $\bm{r}^{*}$ that satisfies $\textsc{TopKMask}(\bm{r}^{*},K)=\textsc{TopKMask}(\mathbb{E}_{p_{\phi}}[\bm{y}],K)$ can be an argmin. One valid solution here is the per-site mean $\bm{\bar{r}}(\phi)=\mathbb{E}_{p_{\phi}}[\bm{y}]$. However, this solution is optimal for a potentially quite loose bound on BPR that approximates the denominator with the constant $K\cdot U$.
Tighter bound suggests the ratio estimator for ranking. Instead of bounding with constant $K\cdot U$, we can upper bound the denominator in Eq. (4) by summing over all $S$ terms instead of the top $K$: $\sum_{s\in\textsc{TopKIDs}(\bm{y})}y_{s}\leq\sum_{s=1}^{S}y_{s}=\bm{1}\cdot\bm{y}$. This bound is tighter when the sum of all entries is less than $K\cdot U$, which is typically true of sparse $\bm{y}$ vectors in our applications. Our how-to-rank problem then becomes
@@ 4 Methods for Training
We now consider various ways to train the parameters of our probabilistic model on a training set that covers $T$ distinct time periods indexed by $t$.
@@ 4.1 Maximum likelihood (ML) estimation
If the model is well-specified and training set size $T$ is large enough, this is a reliable strategy to estimate $\phi$. We could then use the ranking methods from Sec. 3 for where-to-intervene decisions. However, popular wisdom reminds us that “all models are wrong” in some way for real-world data. As we will show in later experiments, fitting a misspecified model via ML estimation can produce $\phi$ that deliver suboptimal BPR, even using the optimal ranking for that $\phi$.
@@ 4.2 Direct loss minimization for BPR
Inspired by the broad goal of direct loss minimization (weiDirectLoss2021), another approach would be to find parameters that minimize our BPR-specific decision making loss $L^{\text{BPR}}$. In this strategy, we seek $\phi$ values that minimize
@@
To train with modern gradient methods, we’d need to compute the gradient $\nabla_{\phi}\mathcal{J}=\sum_{t}\nabla_{\phi}\bm{r}_{t}\cdot\nabla_{r_{t}}L_{t}$. However, technical difficulties arise with each term in this chain rule expansion. Below, we propose practical estimators for each term that overcome these difficulties.
Gradient $\nabla_{\phi}\bm{r}_{t}$. The difficulty here is differentiating through the expectation in Eq. (7), especially when $\bm{y}$ is a discrete random variable (integer counts in our later overdose or wildlife applications). We use the score function trick (mohamed2020MonteCarloGradient), popularized by ranganath2014BBVI yet dating back decades (kleijnen1996OptimizationScoreFunction), sometimes also called REINFORCE (williams1992reinforce). We can draw $M$ samples $\bm{y}^{(m)}_{t}\sim p_{\phi}$, then compute
@@
A downside of this estimator is high variance. We mitigate this with large $M$ values, though future work may use control variates (ranganath2014BBVI; mohamed2020MonteCarloGradient) or try other estimators for gradients of discrete expectations (maddison2017Concrete; dimitriev2021CARMS).
Gradient $\nabla_{r_{t}}L_{t}$. For losses $L$ defined in terms of TopKMask binary vectors, like BPR, it is difficult to compute useful gradients because this loss is flat almost everywhere with respect to the input rankings $\bm{r}$. To overcome this barrier, we leverage recent advances in perturbed optimization (berthetLearningDifferentiablePerturbed2020), also referred to as stochastic smoothing (abernethyPerturbationTechniquesOnline2016). A recent computer vision method (cordonnierDifferentiablePatchSelection2021) shows how these ideas enable selecting a top-K set of patches from a high-resolution image for downstream prediction. We adapt this top-K approach to spatiotemporal forecasting for intervention.
Concretely, cordonnierDifferentiablePatchSelection2021 obtain tractable $J$-sample Monte Carlo estimates of both the top-K indicator vector $\bm{b}=\textsc{TopKMask}(\bm{r},K)$ and the Jacobian $\nabla_{\bm{r}}\bm{b}$ needed for backpropagation. First, we draw $J$ independent samples of a standard Gaussian noise vector of size $S$: $\bm{z}_{j}\sim\mathcal{N}(0,I_{S})$. Then, we compute
@@
The Jacobian $\nabla_{\bm{r}}\bm{\hat{b}}$ is an $S\times S$ matrix, where entry $j,k$ gives the scalar derivative $\frac{\partial\bm{b}_{j}}{\partial r_{k}}$. The conceptual justification for the Jacobian estimator comes from abernethyPerturbationTechniquesOnline2016 and berthetLearningDifferentiablePerturbed2020. Noise level $\sigma>0$ is a hyperparameter that sets the strength of stochastic smoothing. It must be carefully selected in practice to add enough noise so that indicators $\bm{b}_{j}$ change for different samples $z_{j}$, but not too much noise so the $\bm{b}_{j}$ preserve the signal in $\bm{r}$.
@@
We assume the true data-generating distribution is unchanged across train and test time periods. If this does not hold, objectives that just average over $t$ as in Eq. (9) may have disadvantages. Instead we could upweight later $t$, or minimize out-of-sample error as in gupta_decision-aware_2024.
Other methods for decision-aware training. Our approach to direct BPR optimization here is an example of decision-aware or decision-focused training. In the taxonomy of mandi_decision-focused_2024, our approach is in the family of differentiable perturbed optimizers. Other work instead pursues surrogate losses. The SPO+ method (elmachtoubSmartPredictThen2022) finds a convex surrogate for the “smart predict then optimize” optimization problem. The perturbed gradient (PG) method (huang_decision-focused_2024) develops a more sophisticated surrogate, with theory and experiments suggesting utility even with misspecified models. In both cases, surrogate bounds make SGD-based learning tractable for a wide set of optimization tasks, including our knapsack-like BPR problem but also other tasks like shortest path finding.
Downsides of only fitting BPR. When models are misspecified, directly estimating $\phi$ to minimize $\mathcal{J}^{\text{BPR}}$ should yield better BPR than some $\phi^{\prime}$ fit via conventional loss $\mathcal{J}^{\text{NLL}}$, and better BPR means better decisions. However, probabilistic forecasts of near-future outcomes $\bm{y}_{t+1}$ produced by BPR-trained $\phi$ have questionable utility. Nothing in the $\mathcal{J}^{\text{BPR}}$ objective makes $p_{\phi}$ accurately reconstruct even the train set $\bm{y}_{1:T}$; only relative ranking of sites matters. Even with high-quality decisions, a model which produces unlikely forecasts may be difficult to interpret or verify.